# From Efficiency to Adaptivity: A Deeper Look at Adaptive Reasoning in Large Language Models

### v1

@@ 3.1.2 Supervised Fine-tuning and Distillation Methods
A complementary line of work leverages supervised fine-tuning (SFT) or distillation to teach models to produce shorter yet equally valid reasoning chains. C3oT (kang2025c3ot) jointly trains on long–short pairs of chain-of-thought (CoT) exemplars, allowing conditional generation of compact rationales without accuracy loss. CoT-Valve (Ma2025CoTValve) identifies a controllable “length direction” in parameter space and tunes it with paired long/short data for elastic reasoning compression. TokenSkip (Xia2025TokenSkip) prunes semantically redundant tokens and retrains the model under different compression ratios to learn explicit compression control. LS-Mixture SFT (Yu2025LSMixtureSFT) integrates long and short CoT data during fine-tuning, enabling adaptive reasoning depth across instances. Distilling 2-1 (Zhang2024DistillingTwoOne) removes explicit reasoning chains during distillation while retaining correctness, encouraging implicit short reasoning behaviors. SoftCoT (Xu2025SoftCoT) and SoftCoT++ (Xu2025SoftCoTpp) represent thought processes as continuous latent vectors, allowing smooth trajectory distillation that generalizes across prompts. Skip-Thinking (chen2025skipthinkingchunkwisechainofthoughtdistillation) conducts chunk-wise distillation to mitigate gradient saturation in long CoT training. Mixture of Reasonings (MoR) (xiong2025mixturereasoningsteachlarge) jointly trains LLMs on multiple reasoning strategies (e.g., logical, commonsense, numerical) through multi-style CoT supervision, enabling implicit adaptive reasoning style selection without explicit routing. TALE (Han2025TALE) performs token-budget–aware prompt tuning that learns to predict minimal feasible token counts per instance. Length-Instruction Fine-Tuning with Chain-of-Thought (LIFT-COT) (electronics14081662) extends standard supervision with explicit length labels, enabling the model to adhere to user-specified reasoning length or latency constraints while maintaining coherence in its generated rationales. SCoT (wang2025efficientreasoningllmsspeculative) introduces a speculative chain-of-thought framework where a small draft model rapidly generates multiple reasoning trajectories, and a large target model selects or corrects them, achieving up to 3× faster inference with near-target-level accuracy. Together, these supervised and distillation-based methods internalize compression control within model parameters, achieving adaptive reasoning efficiency without reinforcement objectives.
@@ 3.1.3 Learned Controllers and Routers
Another line of work introduces auxiliary controllers or routing modules trained jointly or sequentially with the base model to decide which reasoning path or sub-model to activate. RouteLLM (ong2024routellm) learns a lightweight router from preference data to dispatch easy inputs to cheaper small models and difficult ones to larger models, reducing total compute without degrading accuracy. SelfBudgeter (Li2025SelfBudgeter) trains a two-stage system that first predicts the required reasoning budget and then enforces it through gradient-penalized reinforcement optimization. ThinkSwitcher (Liang2025ThinkSwitcher) extends this idea by training a supervised switching module that decides between concise and elaborate reasoning trajectories based on predicted task difficulty. In Mixture-of-Experts (MoE) models, AdaMOE (zeng2024adamoe) achieves token-adaptive computation by introducing "null experts" that consume zero FLOPs. A load-balancing loss is used during training to teach the model’s router to adaptively send tokens to either true experts for computation or null experts to skip computation, thereby dynamically allocating resources at the token level. Route-To-Reason (RTR) (pan2025routereasonadaptiverouting) learns to predict accuracy–cost trade-offs across model–strategy pairs and dynamically routes each query to the optimal expert combination under a compute budget. Collectively, these works embed adaptivity at the system-decision level and bridge symbolic routing with learned compute control.
@@ 3.2 Training-free Adaptive Reasoning
Training-free adaptive reasoning denotes inference-time mechanisms that modulate the depth, breadth, or structure of reasoning without any parameter update. Formally, a frozen model $f_{\theta}$ generates a reasoning trajectory $r=(t_{1},\dots,t_{n})$ governed by a control policy $\pi(x,t_{i})$ that dynamically decides continuation or termination based on input difficulty or internal uncertainty. Unlike static efficiency control, these approaches allocate computation adaptively for each instance, emphasizing resource-aware and self-regulated inference (Sui2025StopOverthinking; alomrani2025reasoningbudgetsurveyadaptive; zhu2025conciseadaptivethinkinglarge).
@@ 3.2.1 Prompt-conditioned Methods
Explicit token- and instruction-level control. Prompt-conditioned methods introduce explicit instructions, control tokens, or concise templates that determine reasoning length at test time. They embody external adaptivity—computation is modulated through textual conditions rather than learned signals. Lee2025TokenComplexity quantifies token complexity to estimate minimal reasoning spans per instance. Contraied-CoT (Nayab2025ConciseThoughts) employs templates that prioritize conciseness, and enforce concise reasoning by setting explicit length constraints (e.g., ’100 words’). Sketch-of-Thought (aytes2025sketchofthoughtefficientllmreasoning) introduces cognitive-inspired sketching to generate compact, structured reasoning trajectories. Chain-of-Draft (CoD) (Xu2025ChainOfDraft) introduces a “draft-first” reasoning style: the model writes short, essential reasoning steps (at most  5 words) per turn, thus minimizing verbosity while preserving correctness. ConciseHint (Tang2025ConciseHint) integrates continuous hints or contextual anchors for difficulty-aware reasoning. Instruction-level prompt schemes such as Plan-and-Budget (Lin2025PlanBudget) introduce explicit planning and budgeting instructions during inference, enabling the model to allocate reasoning depth and token usage adaptively across sub-questions.
@@ 3.2.2 Feedback-driven Methods
Uncertainty- and entropy-based halting. Feedback-driven adaptation relies on internal feedback such as entropy, confidence, or semantic closure to determine when reasoning should stop or expand. laaouach2025halt defines entropy-threshold halting, where decoding terminates once uncertainty drops below a pre-defined threshold. Yang2025DEER introduces a feedback-driven early exit mechanism that monitors token-level confidence during reasoning to adaptively halt generation once sufficient certainty is reached. mao2025earlystoppingchainofthoughtslarge employs a run-jump test to detect answer convergence, terminating the reasoning process when a sufficient number of consecutive steps yield identical intermediate answers. EAGER (Scalena2025EAGER) leverages token-level entropy to branch into multiple reasoning paths only when high uncertainty is detected. Self-Certainty (Kang2025SelfCertainty) proposes a metric that evaluates answer quality from the model’s own output probabilities, offering a reward-model-free substitute for Best-of-N selection. Sequence-level entropy (zhu2025uncertainty; sharma2025think) and confidence-based filtering methods such as DeepConf(Fu2025DeepConf) further refine this paradigm.
@@
Verifier- and constraint-guided feedback. Beyond internal entropy and confidence, some approaches externalize feedback via verifiers or structural constraints. liu2025answerconvergence detect answer convergence across partial chains and stop early when consistency stabilizes. wang2025nowait prune filler thinking tokens by suppressing keywords (e.g., “Wait”, “Hmm”) during decoding. Speculative or constraint-based schemes such as Speculative Search (wang2025specreasoning) and Value-Guided Search (wang2025valueguided) use acceptance thresholds or value models to admit only high-quality thoughts under limited budgets.
@@
MetaReasoner (sui2025metareasonerdynamicguidanceoptimized) employs a contextual bandit controller that dynamically adjusts reasoning depth during inference based on external validation signals, exemplifying feedback-driven adaptation without reinforcement fine-tuning. These methods remain feedback-driven in essence: reasoning continues or halts based on real-time validation rather than fixed templates.
@@
Consistency- and sampling-based adaptation. A parallel line of work focuses on reasoning breadth—how many rationales to sample before aggregation. Adaptive-Consistency (aggarwal2023letssamplestepstep) and RASC (wan2025reasoningawareselfconsistencyleveraging) estimate consensus and halt when majority agreement stabilizes.
@@
Best-of-NN optimization is improved by Fast BoN (Sun2024FastBoN), which employs efficient sampling algorithms to generate high-scoring responses,and ST-BoN (Wang2025STBoN), which leverages early sampling consistency to identify and truncate suboptimal reasoning paths before full generation, while DPTS (Ding2025DPTS) adaptively prunes search branches. Correctness-first sampling (Li2025SampleSmartNotHard), also known as Sample Smart, further balance accuracy and latency, refining adaptive breadth control for reasoning LLMs.
@@ 3.2.3 Modular Methods
Although implemented through decoding-level control rather than explicit architectural modules, AlphaOne (Zhang2025AlphaOne) employs pacing tokens to regulate cognitive tempo across short-, long-, and tool-augmented reasoning phases. This design mimics modular behavior by dynamically allocating computation among distinct reasoning modes, bridging concise and deliberative thinking within a unified system. CoThink (fan2025cothink) introduces a training-free modular pipeline where an instruct model first drafts a concise reasoning outline, and a thinking model expands it, reducing token usage while maintaining accuracy. Industrial frameworks such as Bag-of-Tricks for Inference-Time Compute (liu2025bagtricksinferencetimecomputation) systematically benchmarks training-free modular techniques that combine generation, verification, and confidence-weighted selection to improve inference-time reasoning efficiency. These systems demonstrate how architectural modularity induces reasoning flexibility without retraining.
@@
A complementary branch combines pre-trained reasoning modules by parameter interpolation or selective ensembling. Long-to-short reasoning via model merging(L2S merging) (Wu2025L2S) reduces average reasoning length while preserving accuracy, and activation-informed merging(ACM)(Yao2025ACM) learns layer-wise coefficients from activation mutual information. For language-specific settings, model merging can transfer reasoning ability from a strong “reasoner” into a target-language model without retraining (Pipatanakul2025LangSpecificMerge). Commercial deployments such as Kimi k1.5 also report hybrid short/long reasoning variants obtained by merging long-CoT and short-CoT specialists (TeamKimi2025Kimi15). These practices share the same anchor that adaptive compute distribution across fixed modules via weight interpolation or expert ensembling.

### v2

@@ 3.1.2 Supervised Fine-tuning and Distillation Methods
A complementary line of work leverages supervised fine-tuning (SFT) or distillation to teach models to produce shorter yet equally valid reasoning chains. C3oT (Kang et al., 2025a) jointly trains on long–short pairs of chain-of-thought (CoT) exemplars, allowing conditional generation of compact rationales without accuracy loss. CoT-Valve (Ma et al., 2025) identifies a controllable “length direction” in parameter space and tunes it with paired long/short data for elastic reasoning compression. TokenSkip (Xia et al., 2025a) prunes semantically redundant tokens and retrains the model under different compression ratios to learn explicit compression control. LS-Mixture SFT (Yu et al., 2025a) integrates long and short CoT data during fine-tuning, enabling adaptive reasoning depth across instances. Distilling 2-1 (Yu et al., 2024) removes explicit reasoning chains during distillation while retaining correctness, encouraging implicit short reasoning behaviors. SoftCoT (Xu et al., 2025b) and SoftCoT++ (Xu et al., 2025c) represent thought processes as continuous latent vectors, allowing smooth trajectory distillation that generalizes across prompts. Skip-Thinking (Chen et al., 2025b) conducts chunk-wise distillation to mitigate gradient saturation in long CoT training. Mixture of Reasonings (MoR) (Xiong et al., 2025) jointly trains LLMs on multiple reasoning strategies (e.g., logical, commonsense, numerical) through multi-style CoT supervision, enabling implicit adaptive reasoning style selection without explicit routing. TALE (Han et al., 2025) performs token-budget–aware prompt tuning that learns to predict minimal feasible token counts per instance. Length-Instruction Fine-Tuning with Chain-of-Thought (LIFT-COT) (Chen and Li, 2025) extends standard supervision with explicit length labels, enabling the model to adhere to user-specified reasoning length or latency constraints while maintaining coherence in its generated rationales. SCoT (Wang et al., 2025b) introduces a speculative chain-of-thought framework where a small draft model rapidly generates multiple reasoning trajectories, and a large target model selects or corrects them, achieving up to 3× faster inference with near-target-level accuracy. Together, these supervised and distillation-based methods internalize compression control within model parameters, achieving adaptive reasoning efficiency without reinforcement objectives.
@@ 3.1.3 Learned Controllers and Routers
Another line of work introduces auxiliary controllers or routing modules trained jointly or sequentially with the base model to decide which reasoning path or sub-model to activate. RouteLLM (Ong et al., 2025) learns a lightweight router from preference data to dispatch easy inputs to cheaper small models and difficult ones to larger models, reducing total compute without degrading accuracy. SelfBudgeter (Li et al., 2025b) trains a two-stage system that first predicts the required reasoning budget and then enforces it through gradient-penalized reinforcement optimization. ThinkSwitcher (Liang et al., 2025) extends this idea by training a supervised switching module that decides between concise and elaborate reasoning trajectories based on predicted task difficulty. In Mixture-of-Experts (MoE) models, AdaMOE (Zeng et al., 2024) achieves token-adaptive computation by introducing "null experts" that consume zero FLOPs. A load-balancing loss is used during training to teach the model’s router to adaptively send tokens to either true experts for computation or null experts to skip computation, thereby dynamically allocating resources at the token level. Route-To-Reason (RTR) (Pan et al., 2025) learns to predict accuracy–cost trade-offs across model–strategy pairs and dynamically routes each query to the optimal expert combination under a compute budget. Collectively, these works embed adaptivity at the system-decision level and bridge symbolic routing with learned compute control.
@@ 3.2 Training-free Adaptive Reasoning
Training-free adaptive reasoning denotes inference-time mechanisms that modulate the depth, breadth, or structure of reasoning without any parameter update. Formally, a frozen model $f_{\theta}$ generates a reasoning trajectory $r=(t_{1},\dots,t_{n})$ governed by a control policy $\pi(x,t_{i})$ that dynamically decides continuation or termination based on input difficulty or internal uncertainty. Unlike static efficiency control, these approaches allocate computation adaptively for each instance, emphasizing resource-aware and self-regulated inference (Sui et al., 2025a; Alomrani et al., 2025; Zhu and Li, 2025).
@@ 3.2.1 Prompt-conditioned Methods
Explicit token- and instruction-level control. Prompt-conditioned methods introduce explicit instructions, control tokens, or concise templates that determine reasoning length at test time. They embody external adaptivity—computation is modulated through textual conditions rather than learned signals (Cai et al., 2025). Lee et al. (2025) quantifies token complexity to estimate minimal reasoning spans per instance. Contraied-CoT (Nayab et al., 2025) employs templates that prioritize conciseness, and enforce concise reasoning by setting explicit length constraints (e.g., ’100 words’). Sketch-of-Thought (Aytes et al., 2025) introduces cognitive-inspired sketching to generate compact, structured reasoning trajectories. Chain-of-Draft (CoD) (Xu et al., 2025a) introduces a “draft-first” reasoning style: the model writes short, essential reasoning steps (at most  5 words) per turn, thus minimizing verbosity while preserving correctness. ConciseHint (Tang et al., 2025) integrates continuous hints or contextual anchors for difficulty-aware reasoning. Instruction-level prompt schemes such as Plan-and-Budget (Lin et al., 2025) introduce explicit planning and budgeting instructions during inference, enabling the model to allocate reasoning depth and token usage adaptively across sub-questions.
@@ 3.2.2 Feedback-driven Methods
Uncertainty- and entropy-based halting. Feedback-driven adaptation relies on internal feedback such as entropy, confidence, or semantic closure to determine when reasoning should stop or expand. Laaouach (2025) defines entropy-threshold halting, where decoding terminates once uncertainty drops below a pre-defined threshold. Yang et al. (2025) introduces a feedback-driven early exit mechanism that monitors token-level confidence during reasoning to adaptively halt generation once sufficient certainty is reached. Mao et al. (2025) employs a run-jump test to detect answer convergence, terminating the reasoning process when a sufficient number of consecutive steps yield identical intermediate answers. EAGER (Scalena et al., 2025) leverages token-level entropy to branch into multiple reasoning paths only when high uncertainty is detected. Self-Certainty (Kang et al., 2025b) proposes a metric that evaluates answer quality from the model’s own output probabilities, offering a reward-model-free substitute for Best-of-N selection. Sequence-level entropy (Zhu et al., 2025; Sharma and Chopra, 2025) and confidence-based filtering methods such as DeepConf(Fu et al., 2025) further refine this paradigm.
@@
Verifier- and constraint-guided feedback. Beyond internal entropy and confidence, some approaches externalize feedback via verifiers or structural constraints. Liu and Wang (2025) detect answer convergence across partial chains and stop early when consistency stabilizes. Wang et al. (2025a) prune filler thinking tokens by suppressing keywords (e.g., “Wait”, “Hmm”) during decoding. Speculative or constraint-based schemes such as Speculative Search (Wang et al., 2025f) and Value-Guided Search (Wang et al., 2025c) use acceptance thresholds or value models to admit only high-quality thoughts under limited budgets.
@@
MetaReasoner (Sui et al., 2025b) employs a contextual bandit controller that dynamically adjusts reasoning depth during inference based on external validation signals, exemplifying feedback-driven adaptation without reinforcement fine-tuning. These methods remain feedback-driven in essence: reasoning continues or halts based on real-time validation rather than fixed templates.
@@
Consistency- and sampling-based adaptation. A parallel line of work focuses on reasoning breadth—how many rationales to sample before aggregation. Adaptive-Consistency (Aggarwal et al., 2023) and RASC (Wan et al., 2025) estimate consensus and halt when majority agreement stabilizes.
@@
Best-of-NN optimization is improved by Fast BoN (Sun et al., 2024), which employs efficient sampling algorithms to generate high-scoring responses,and ST-BoN (Wang et al., 2025e), which leverages early sampling consistency to identify and truncate suboptimal reasoning paths before full generation, while DPTS (Ding et al., 2025) adaptively prunes search branches. Correctness-first sampling (Li et al., 2025a), also known as Sample Smart, further balance accuracy and latency, refining adaptive breadth control for reasoning LLMs.
@@ 3.2.3 Modular Methods
Although implemented through decoding-level control rather than explicit architectural modules, AlphaOne (Zhang et al., 2025) employs pacing tokens to regulate cognitive tempo across short-, long-, and tool-augmented reasoning phases. This design mimics modular behavior by dynamically allocating computation among distinct reasoning modes, bridging concise and deliberative thinking within a unified system. CoThink (Fan et al., 2025) introduces a training-free modular pipeline where an instruct model first drafts a concise reasoning outline, and a thinking model expands it, reducing token usage while maintaining accuracy. Industrial frameworks such as Bag-of-Tricks for Inference-Time Compute (Liu et al., 2025a) systematically benchmarks training-free modular techniques that combine generation, verification, and confidence-weighted selection to improve inference-time reasoning efficiency. These systems demonstrate how architectural modularity induces reasoning flexibility without retraining.
@@
A complementary branch combines pre-trained reasoning modules by parameter interpolation or selective ensembling. Long-to-short reasoning via model merging(L2S merging) (Wu et al., 2025) reduces average reasoning length while preserving accuracy, and activation-informed merging(ACM)(Yao et al., 2025) learns layer-wise coefficients from activation mutual information. For language-specific settings, model merging can transfer reasoning ability from a strong “reasoner” into a target-language model without retraining (Pipatanakul et al., 2025). Commercial deployments such as Kimi k1.5 also report hybrid short/long reasoning variants obtained by merging long-CoT and short-CoT specialists (Kimi, 2025). These practices share the same anchor that adaptive compute distribution across fixed modules via weight interpolation or expert ensembling.